OpenStudy (samigupta8):

If f(x) is continuous and f(9/2)=2/9 then Lim f(1-cos3x÷x^2) is equal to... Where x tends to 0

OpenStudy (samigupta8):

@sparrow2

OpenStudy (samigupta8):

@faiqraees

OpenStudy (sparrow2):

if the function f(x)=1/x it seems so

OpenStudy (samigupta8):

How can we say that ?? I mean it can be anything ....how can u arrive at it so easily just by luking at one value of f(x) we can't do injustice to whole function....

OpenStudy (sparrow2):

that's very true.that why i wrote it "seems" .we can't judge that

OpenStudy (samigupta8):

Okk sorry for that..

OpenStudy (samigupta8):

Guys....the division sign is under the whole expression and not just cos3x

OpenStudy (mayankdevnani):

i got 2/9 !

OpenStudy (samigupta8):

How???

OpenStudy (mayankdevnani):

is it right or wrong ?

OpenStudy (samigupta8):

Ryt.....yrr mayank kse kiya tumne abb yeh

OpenStudy (mayankdevnani):

lol tukka

OpenStudy (samigupta8):

Pitoge shi tarika btao

OpenStudy (mayankdevnani):

nahi aata

OpenStudy (samigupta8):

Pls....do tell mayank

OpenStudy (mayankdevnani):

okay

OpenStudy (mayankdevnani):

so can i type whole solution ?

OpenStudy (samigupta8):

Yep

OpenStudy (mayankdevnani):

\[\large \bf 1-\cos 3x=2 \sin^2(\frac{3x}{2})\] \[\large \bf \lim_{x \rightarrow 0}f(\frac{2 \sin^2(\frac{3x}{2})}{x^2})\]